# Sat-DN: Implicit Surface Reconstruction from Multi-View Satellite Images with Depth and Normal Supervision

### v1

@@ V-E2 Depth Regularization
Notably, due to imbalances in lighting within the scene, errors in surface position estimation occurred when reconstructing building rooftops, necessitating the subsequent introduction of normal constraints. Fig. 9 lists the number of retained spatial 3D points and their reprojection errors in 2D images after bundle adjustment for each scene. It can be observed that the number of spatial points and their reprojection errors impact the reconstruction results of our method. The number of 3D points positively correlates with reconstruction quality, while reprojection error is negatively correlated.
@@ V-E3 Normal Regularization
VI Conclusion

### v2

@@ V-E2 Depth Regularization
Figure 10: Least square depth fusion errors of each scene in DFC2019 dataset.
@@
Notably, due to imbalances in lighting within the scene, errors in surface position estimation occurred when reconstructing building rooftops, necessitating the subsequent introduction of normal constraints. Fig. 9 lists the number of retained spatial 3D points and their reprojection errors in 2D images after bundle adjustment for each scene. It can be observed that the number of spatial points and their reprojection errors impact the reconstruction results of our method. The number of 3D points positively correlates with reconstruction quality, while reprojection error is negatively correlated. In addition, we recorded the final least square losses means and medians of the depth fusion in Fig. 10. In scenarios with large height differences or complicated geometry, such as JAX_214 and OMA_248, although the depth estimation model can obtain the relative elevation, this relative elevation cannot accurately correspond to the real height difference. However, in the vast majority of scenarios, the error of depth fusion is relatively small and can guide the model to converge to the correct geometry, thereby alleviating the ambiguity of optimization brought about by fluctuations in pixel photometric values.
The Depth Anything V2[20] model we use is one of the SOTA models in the computer vision community, which is trained by real-world images. Despite the limitations of using computer vision models in satellite images, the experiments presented above prove that the computer vision model can achieve a reasonable relative depth estimation in satellite images. Since this study focuses on surface reconstruction, the depth estimation content is not ablated.
@@ V-E3 Normal Regularization
VI Discussion
VI-A Comparison between Nerf-based and Traditional Methods
Traditional methods for processing satellite imagery primarily rely on sparse or stereo matching techniques. However, these approaches depend on complete visual correspondences and highly accurate matching, which often result in discontinuities and voids in the reconstructed point clouds and meshes. Such incomplete and fragmented outputs are typically unsuitable for downstream applications, especially in scenarios that demand high-fidelity 3D representations. Moreover, traditional methods are prone to the influence of dynamic objects, often introducing outliers into the reconstruction.
In contrast, NeRF-based methods offer a significant advantage by generating watertight and continuous meshes that are structurally coherent and largely free from such artifacts. These meshes can be directly manipulated using standard 3D modeling software, thereby enhancing their practical applicability.
Although NeRF-based approaches generally require longer processing times, the improved quality and completeness of the reconstructed geometry often justify the additional computational cost. With ongoing advancements in hardware, this gap is expected to gradually diminish.
VII Conclusion